For an account that earns simple annual interest,find the interest and the balance of the account.
P= $4500 ,r = 4%,t = 15 years
Any help with this equation would Greatly appreciated.

Answer by checkley71(8403)

I=P(1+R)^T-P
I=4500(1+.04)^15-4500
I=4500*1.8-4500
I=810-4500
I=3600 IS THE INTEREST.
4500+3600=8100 IS THE BALANCE AT THE END OF 15 YEARS.